OpenStudy (anonymous):

The sum of the squares of two consecutive odd integers is 74. Find the two integers.

OpenStudy (dumbcow):

odd integer can be represented as 2n+1 and the next odd int would 2n+3 so you have: (2n+1)^2 + (2n+3)^2 = 74 solve for n

OpenStudy (anonymous):

i dont understand it..

OpenStudy (anonymous):

i dont remember how to solve n

OpenStudy (dumbcow):

hmm i guess an easier way is to look at perfect squares and which ones add up to 74 1,3 = 1^2 + 3^2 = 10 3,5 = 3^2 + 5^2 = 34 ....

OpenStudy (anonymous):

so its 3 and 5?

OpenStudy (dumbcow):

just remember when you find it that their negative counterparts are answers too because we square them -5,-3, = (-5)^2 + (-3)^2 = 34 ...

OpenStudy (dumbcow):

no sum of 3 and 5 squared is 34...keep going try 5,7 7,9 9,11 has to equal 74

OpenStudy (anonymous):

5 and 7

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

Your first method was nicer :( You can't brute force it once you get up to bigboy numbers